### v4

# An entanglement witness and an effective criterion for bipartite states of different dimensions

Abstract
We present a simple approach for generation of a diverse set of positive maps between spaces of different dimensions. The proposed method enables the construction of Entanglement Witnesses tailored for systems in $d_{1}\times d_{2}$ dimensions. We then show how application of this entanglement witness to a general density matrix leads to a simple calculable criterion in closed form. As a support for the power of this criterion, we apply it to a variety of (one-parameter) PPT entangled states and show that the ranges of the parameters where these states are entangled are much larger than previously reported.
Introduction
Entanglement stands out as one of the most remarkable distinctions between the quantum and classical realms, playing a pivotal role in numerous quantum protocols, for review papers see [1, 2, 3]. As a result, the creation, manipulation, application, and identification of entanglement have profound significance. The identification of entangled states, in particular, presents a formidable challenge, underscored by its classification as an NP-Hard problem in mathematical terms [4]. Hence, the pursuit of viable techniques for detecting entangled states takes on paramount importance.
Entangled states are referred to as states that are not separable, and separable states refer to states that can be written as follows[5]:
$\rho=\sum p_{i}\rho_{i}^{(1)}\otimes\rho_{i}^{(2)},$ (1)
where $0\leq p_{i}\leq 1$ and $\sum p_{i}=1$. One strategy for identifying entanglement involves investigating whether a state can be expressed in the aforementioned form. This, however, poses a formidable challenge. As a result, there is a pressing need for more effective methods for the detection of entanglement. Over time, diverse techniques have emerged for this purpose. The Peres criterion is considered one of the primary methods in this regard[6]: It’s relatively straightforward to see that the partial transpose of all separable states is positive. This is seen simply by noting that for any state of the form (1), one has
$(I\otimes T)(\rho)\geq 0,$ (2)
where $T$ is the transpose, hence the name partial transpose. Therefore, if the partial transpose of a quantum state is not positive, it must be entangled. However, this criterion alone cannot identify all types of quantum entangled states because there are entangled states known as PPT, whose partial transpose are positive. In fact separable states are a subset of PPT states and only when $d_{1}d_{2}\leq 6$, these two sets are equal. Here $d_{1}$ and $d_{2}$ are the dimensions of the two quantum systems. The effectiveness of the Peres criterion (2) hinges on a basic property of the transpose map $T$ which is a positive and yet not completely positive map. Therefore it was natural to search for more general types of positive maps lacking the property of complete positivity and thus constructing more general Entanglement Witnesses than the simple partial transpose. In fact, it has been known that any entanglement witness $W$ can be written in the form
$W(M)=(I\otimes M)|\phi^{+}\rangle\langle\phi^{+}|,$
where $M$ is a positive map and $|\phi^{+}\rangle=\frac{1}{\sqrt{d}}\sum_{\mu=0}^{d-1}|\mu\mu\rangle$, is a maximally entangled state[7]. In other words, $W(M)$ is the Choi-matrix of the map $M$, which due to its lack of complete positivity can have negative eigenvalue. Such a witness has the property
$\operatorname{Tr}(W\sigma)\geq 0,\quad\forall\sigma\in S,$
where $S$ is the set of all separable states. Alternatively, one can think of $W$ as a block-positive operator, i.e.
$\langle\psi,\phi|W|\psi,\phi\rangle,\ \forall\ \psi\in H_{1}\ ,\forall\ \phi\in H_{2},$
which has negative eigenvalue when acting on the whole Hilbert space $H_{1}\otimes H_{2}$. Thus the problem of constructing entanglement witnesses becomes equivalent to constructing positive but not completely positive maps. Based on this equivalence, entanglement witnesses were introduced in many works, see [7, 8, 9, 10] and references therein. For more recent works, see [11, 12, 13]. In practical situations, where the two particles are far apart in remote laboratories, application of any entanglement witnesses found so far, requires a common reference frame between the two parties holding the particles. It was recently shown in [14] that by a process called incomplete teleportation, all these entangled states are still able to detect entanglement of remote parties, even in the absence of a common frame of reference.
In contrast with completely positive maps which due to the Kraus theorem can be easily constructed, construction of positive maps is not a straightforward task, and appropriate methods need to be proposed for this purpose. In addition to the finding new classes of witnesses, the relation of EW’s and positive maps, have led to many new insights on both [8]. For example, a basic distinction refers to decomposable witnesses which are of the form
$W=P+Q^{\Gamma},$
where $P,Q$ are positive operators and $\Gamma$ is the operator of partial transposition. Any witness that cannot be written in the above form is called indecomposable. It can be proved that decomposable witnesses cannot identify PPT entangled states. In this sense these are somewhat weaker than indecomposable witnesses. One can also characterize witnesses by their optimality and extremality [8], a subject which we will not deal with in this work. The line that we follow is inspired by a number of recent works [15, 16, 17, 18, 19, 20] in which specific types of positive maps have been defined and the corresponding witnesses have been constructed.
It is worth noting that in almost all the mentioned articles, the entanglement witnesses constructed have been designed for two systems of equal dimensions. In this paper, we propose a simple method for identification of entanglement in systems with dimensions of $d_{1}\times d_{2}$. This dimension-specific applicability enhances the versatility of the method, making it a valuable tool in different practical scenarios. Often, due to practical constraints in laboratory setups, certain measurements may be more accessible or accurate than others. Our method elegantly accommodates this by allowing for the construction of entanglement witnesses based solely on those restricted measurements which are feasible in the given experimental context(See examples).
The structure of this paper is as follows: In Section 1, we examine the main idea of this article, which is to construct a new class of positive maps. In Section 2, an alternative method is utilized to obtain the same results that presented in Section 1. In section 3 we provide a simple criterion built with our Entanglement Witness which easily detects entangled states. In Section 4, we present several examples which shows the power of our criterion in detecting PPT entangled states. We end up the paper with a conclusion.
1 Construction of positive maps for arbitrary different dimensions
Consider the Hilbert space $H_{d_{1}}$ of dimension $d_{1}$ and let a pure state $X\in D(H_{d_{1}})$, (the space of states on $H_{d_{1}}$) be written as
$X=\frac{1}{d_{1}}I_{d_{1}}+\sum_{i=1}^{d_{1}^{2}-1}x_{i}\Gamma_{i},$
where $\{\Gamma_{i},\ \ \ i=1\cdots d_{1}^{2}-1\}$ are a set of orthonormal traceless Hermitian operators $\in L(H_{d_{1}})$, i.e.
$\operatorname{Tr}(\Gamma_{i}\Gamma_{j})=\delta_{ij},\ \ \ \operatorname{Tr}\Gamma_{i}=0.$
Purity of the state $X$ (i.e. $\operatorname{Tr}X^{2}=1$) and the above normalization conditions restricts the norm of the vector ${\bf x}=(x_{1},\cdots x_{d_{1}^{2}-1})$ to
${\bf x}\cdot{\bf x}=\frac{d_{1}-1}{d_{1}},$ (3)
where ${\bf x}\cdot{\bf x}:=\sum_{i=1}^{d_{1}^{2}-1}x_{i}^{2}$.
Let $\Phi:D(H_{d_{1}})\longrightarrow D(H_{d_{2}})$ be unital and trace-preserving map which acts as follows:
$\Phi(\frac{1}{d_{1}}I_{d_{1}})=\frac{1}{d_{2}}I_{d_{2}}\ \ \ \ ,\ \ \ \Phi(\Gamma_{i})=\frac{1}{d_{2}}\sum_{k=1}^{d_{2}^{2}-1}\Lambda_{ik}\Omega_{k}.$ (4)
In fact, $\Lambda_{ik}$ are real numbers representing the characterization of the map $\Phi$. Here, our definition remains consistent with the trace-preserving property of the map, in contrast to the definition presented in [18], which is grounded in the property of such a map that takes an identity matrix in $D(H_{d_{1}})$ to a identity matrix in $D(H_{d_{2}})$. The factor $\frac{1}{d_{2}}$ in the second equation is written for later convenience, and $\{\Omega_{k},\ \ \ k=1\cdots d_{2}^{2}-1\}$ are a set of orthonormal traceless Hermitian operators $\in L(H_{d_{2}})$, i.e.
$\operatorname{Tr}(\Omega_{k}\Omega_{l})=\delta_{kl},\ \ \ \operatorname{Tr}\Omega_{k}=0.$
Then $\Phi$ maps $X$ to
$\Phi(X)=\frac{1}{d_{2}}I_{d_{2}}+\frac{1}{d_{2}}\sum_{i=1}^{d_{1}^{2}-1}\sum_{k=1}^{d_{2}^{2}-1}x_{i}\Lambda_{ik}\Omega_{k}.$ (5)
Remark 1
Up to now we have assumed that $X$ is a pure state. This can be extended to arbitrary operatores $X\in L(H_{d_{1}})$ as follows:
$\Phi(X)=\frac{\operatorname{Tr}(X)}{d_{2}}I_{d_{2}}+\frac{1}{d_{2}}\sum_{i=1}^{d_{1}^{2}-1}\sum_{k=1}^{d_{2}^{2}-1}x_{i}\Lambda_{ik}\Omega_{k}.$
To ensure positivity of the map $\Phi$, we now use Mehta’s Lemma [21] which we state below:
Mehta’s Lamma (1989): Let $A$ be a Hermitian matrix of dimension $D$, if
$\operatorname{Tr}(A^{2})\leq\frac{(\operatorname{Tr}A)^{2}}{D-1},$ (6)
then $A$ is positive.
In view of linearity of the map $\Phi$, to prove positivity of the map $\Phi$, it is sufficient to prove that $\Phi(X)$ is positive for any pure state $X$. That is the map $\Phi$ is positive if the following condition holds
$\operatorname{Tr}(\Phi(X)^{2})\leq\frac{(\operatorname{Tr}\Phi(X))^{2}}{d_{2}-1}=\frac{1}{d_{2}-1}.$ (7)
To this end, we calculate $\operatorname{Tr}(\Phi(X)^{2})$, which in view of orthonormality condition (4), turns out to be
$\operatorname{Tr}(\Phi(X)^{2})=\frac{1}{d_{2}}+\frac{1}{d_{2}^{2}}{\bf x}^{T}{\Lambda}\Lambda^{T}{\bf x},$
where the matrix $\Lambda$ is composed of $\Lambda_{ij}$ elements. In view of (6), we demand that
$\frac{1}{d_{2}}+\frac{1}{d_{2}^{2}}{\bf x}^{T}{\Lambda}\Lambda^{T}{\bf x}\leq\frac{1}{d_{2}-1},\hskip 28.45274pt\forall{\bf x}.$
We can now replace this with an even stronger inequality, namely
$\frac{1}{d_{2}}+\frac{1}{d_{2}^{2}}\lambda_{max}(\Lambda{\Lambda^{T}}){\bf x}\cdot{\bf x}\leq\frac{1}{d_{2}-1},$ (8)
where $\lambda_{max}:=\lambda_{max}(\Lambda\Lambda^{T})$, is the largest eigenvalue of the matrix $\Lambda\Lambda^{T}$, or equivalently $\lambda_{max}$ is the square of the largest singular value of the matrix $\Lambda^{T}$. Inserting (8) in (7) and using (3), yields the following constraint on $\lambda_{max}$
$\lambda_{max}\leq\frac{d_{1}d_{2}}{(d_{1}-1)(d_{2}-1)}.$ (9)
The EW corresponding to the map (5) is given by [7]
$W=(I_{d_{1}}\otimes\Phi)|\beta\rangle\langle\beta|,$
where $|\beta\rangle=\frac{1}{\sqrt{d}}\sum_{\mu=0}^{d_{1}-1}|\mu,\mu\rangle$ is a maximally entangled state. To do this calculation, we extend the orthonormal sets $\{\Gamma_{i}\}$ and $\{\Omega_{k}\}$ respectively to orthonormal bases for the space of matrices $M_{d_{1}}$ and $M_{d_{2}}$, namely to $\{\Gamma_{\mu}\}=\{\Gamma_{0}=\frac{I}{\sqrt{d_{1}}},\Gamma_{i}\}$ and $\{\Omega_{\mu}\}=\{\Omega_{0}=\frac{I}{\sqrt{d_{2}}},\Omega_{i}\}$ and taking remark 1 into account, we write
$W$ $=$ $\frac{1}{d_{1}}\sum_{\mu,\nu}|\mu\rangle\langle\nu|\otimes\Phi(|\mu\rangle\langle\nu|)$ (10)
$=$ $\frac{1}{d_{1}}\sum_{\mu,\nu}|\mu\rangle\langle\nu|\otimes\sum_{\alpha=0}^{d_{1}^{2}-1}\sum_{\beta=0}^{d_{2}^{2}-1}\Lambda_{\beta\alpha}\operatorname{Tr}(|\mu\rangle\langle\nu|\Gamma_{\beta})\Omega_{\alpha}$
$=$ $\frac{1}{d_{1}}\sum_{\alpha,\beta}\Lambda_{\beta\alpha}\Gamma^{T}_{\beta}\otimes\Omega_{\alpha}.$ (11)
where $\Lambda_{00}=\frac{1}{d_{2}},\Lambda_{0i}=0,\Lambda_{i0}=0$. After multiplication by an overall factor $d_{1}d_{2}$, this leads to the following witness
$W=I_{d_{1}}\otimes I_{d_{2}}+\sum_{i=1}^{d_{1}^{2}-1}\sum_{k=1}^{d_{2}^{2}-1}{\Lambda}_{ik}\Gamma^{T}_{i}\otimes\Omega_{k}.$
Since the set $\Gamma^{T}_{i}$ is also an orthonormal set, we can rename it to $\Gamma_{i}$. Thus, the following matrix can be a entanglement witness too:
$W=I_{d_{1}}\otimes I_{d_{2}}+\sum_{i=1}^{d_{1}^{2}-1}\sum_{k=1}^{d_{2}^{2}-1}{\Lambda}_{ik}\Gamma_{i}\otimes\Omega_{k}.$ (12)
Therefore for any two dimensions, any operator of the above form can be an entanglement witness, provided that the condition (9) holds. That condition in fact proves Block positivity of this Witness. Whether or not a $W$ constructed in this way, has also a negative eigenvalue depends on the matrix $\Lambda$ that we choose. In all previous papers that were mentioned in the introduction, $\Lambda$ should be orthogonal. These comprise a much smaller class compared to the $\Lambda$ matrices in this paper. Moreover, finding orthogonal matrices in high dimensions is not a straightforward task. In this paper, one can use any matrix with any properties to construct entanglement witnesses, as long as the largest eigenvalue of $\Lambda\Lambda^{T}$ satisfies eq. (9).
In the next section, we derive the same result of eq. (12) in an alternative way by focusing on the EW itself, rather than the positive map which leads to it via its Choi-matrix. In view of this alternative construction, one can first choose the local operators (according to the experimental restrictions) and then tuning the coefficients for defining the final entanglement witness. This is a significant advantage compared to methods where a positive map may lead to a witness which requires non-feasible measurements.
2 An alternative construction of Entanglement Witnesses
In this section we directly apply Mehta’s Lemma to a form of EW which we take as follows:
$W=I_{d_{1}}\otimes I_{d_{2}}+\sum_{i=1}^{d_{1}^{2}-1}\sum_{j=1}^{d_{2}^{2}-1}\Lambda_{ij}\Gamma_{i}\otimes\Omega_{j}.$
We demand that $W$ be block-positive. This means that for any vector $|v\rangle\in H_{d_{2}}$, the following matrix should be positive
$B\equiv\langle v|W|v\rangle=I_{d_{1}}+\sum_{i=1}^{d_{1}^{2}-1}\sum_{j=1}^{d_{2}^{2}-1}\Lambda_{ij}\Gamma_{i}\langle v|\Omega_{j}|v\rangle.$
Now we use the following:
$\operatorname{Tr}(B)=d_{1},$
and
$\operatorname{Tr}(B^{2})=d_{1}+\sum_{i,j=1}^{d_{2}^{2}-1}(\Lambda^{T}\Lambda)_{ij}\langle v|\Omega_{i}|v\rangle\langle v|\Omega_{j}|v\rangle,$
where we have used the relations $\operatorname{Tr}(\Gamma_{i})=0,\ \ \ \operatorname{Tr}(\Gamma_{i}\Gamma_{j})=\delta_{ij}$. We now resort to the Mehta’s Lemma and demand that the following condition hold for all the vectors $|v\rangle$
$\operatorname{Tr}(B^{2})\leq\frac{\operatorname{Tr}(B)^{2}}{d_{1}-1}.$
This leads to
[…]
Defining the vector ${\bf q}$ with components $q_{i}:=\langle v|\Omega_{i}|v\rangle$, this can be written as
[…]
or in a stronger form
$d_{1}+\lambda_{max}(\Lambda^{T}\Lambda){\bf q}.{\bf q}\leq\frac{d_{1}^{2}}{d_{1}-1}.$ (13)
It remains to calculate ${\bf q}.{\bf q}$, and this is easily found by noting that
${\bf q}.{\bf q}=\sum_{i=1}^{d_{2}^{2}-1}\langle v|\Omega_{i}|v\rangle\langle v|\Omega_{i}|v\rangle=\sum_{i=1}^{d_{2}^{2}-1}\langle v,v|\Omega_{i}\otimes\Omega_{i}|v,v\rangle.$
Using the identity
$\frac{1}{d}I+\sum_{i}\Omega_{i}\otimes\Omega_{i}=P,$ (14)
where $P$ is the permutation operator, we find that
${\bf q}.{\bf q}=\langle v,v|P-\frac{1}{d_{2}}I|v,v\rangle=1-\frac{1}{d_{2}}.$
Inserting this in (13), leads to the final form for the upper bound of $\lambda_{max}$, namely
$\lambda_{max}(\Lambda^{T}\Lambda)\leq\frac{d_{1}d_{2}}{(d_{1}-1)(d_{2}-1)},$ (15)
which is identical to the bound obtained by the positive map in section 1. However we should also ensure that $W$ is not trivially positive, that is we should remove the possibility that all its eigenvalues be positive. We again use Mehta’s lemma. To this end we use
$\operatorname{Tr}(W)=d_{1}d_{2},$
and
$\operatorname{Tr}(W^{2})=d_{1}d_{2}+\operatorname{Tr}(\Lambda^{T}\Lambda),$
and ensure that the following inequality be satisfied
$\operatorname{Tr}(W^{2})>\frac{\operatorname{Tr}(W)^{2}}{d_{1}d_{2}-1},$
or
$\frac{d_{1}d_{2}}{d_{1}d_{2}-1}<\operatorname{Tr}(\Lambda^{T}\Lambda).$ (16)
Combining (15) and (16), we arrive at the final condition for the parameters of the matrix $\Lambda$, in order that $W$ can be an entanglement witness:
$\frac{d_{1}d_{2}}{d_{1}d_{2}-1}<\operatorname{Tr}(\Lambda^{T}\Lambda),\ \ \ \ {\rm and}\ \ \ \lambda_{max}\leq\frac{d_{1}d_{2}}{(d_{1}-1)(d_{2}-1)}.$ (17)
Note that the second condition is sufficient for the block-positivity of $W$, while the first condition is necessary for $W$, ensuring that it is not a positive matrix.
3 A simple criterion
Consider now the form of the witness in equation (12) subject to the condition
$\lambda_{max}\leq\frac{d_{1}d_{2}}{(d_{1}-1)(d_{2}-1)},$ (18)
which is sufficient for its block-positivity. Let an arbitrary bi-partite state be expressed as follows in an orthogonal trace-less basis:
$\rho=\frac{1}{d_{1}d_{2}}I_{d_{1}}\otimes I_{d_{2}}+\sum_{i=1}^{d_{1}^{2}-1}{a}_{i}\Gamma_{i}\otimes I_{d_{2}}+\sum_{j=1}^{d_{2}^{2}-1}{b}_{j}I_{d_{1}}\otimes\Omega_{j}+\sum_{i=1}^{d_{1}^{2}-1}\sum_{j=1}^{d_{2}^{2}-1}{r}_{ij}\Gamma_{i}\otimes\Omega_{j}.$
It is readily seen that
$\operatorname{Tr}(W\rho)=1+\sum_{ij}\Lambda_{ij}r_{ij}=1+\operatorname{Tr}(R\Lambda^{T}),$ (19)
where the matrix $R$ is composed of $r_{ij}$ elements. In order to make $\operatorname{Tr}(W\rho)<0$, which at the same time proves that $W$ is indeed an entanglement witness (i.e. has a negative eigenvalue) and detects the entanglement of $\rho$, we try to choose $\Lambda$ in such a way to minimize $tr(R\Lambda^{T})$, while maintaining the condition (18). To achieve this, we first express the matrix $R$ using singular value decomposition:
$R=USV.$
Here we have used the fact that $R$ is a real matrix, which implies the matrices $U\in M_{d_{1}\times d_{1}}$ and $V\in M_{d_{2}\times d_{2}}$ are real orthogonal matrices. The matrix $S\in M_{d_{1}\times d_{2}}$ is a rectangular diagonal matrix with non-negative entries on the main diagonal. Now we choose the matrix $\Lambda^{T}$ as follows:
$\Lambda^{T}=V^{T}LU^{T},$
where $L\in M_{d_{2}\times d_{1}}$ is chosen to be a rectangular diagonal matrix whose entries on its main diagonal are equal to $l_{i}=-\sqrt{\frac{d_{1}d_{2}}{(d_{1}-1)(d_{2}-1)}}$. This leads to
$\Lambda^{T}\Lambda=V^{T}(LL^{T})V,$
where $LL^{T}\in M_{d_{2}\times d_{2}}$ is a matrix whose diagonal elements are either zero or equal to $\frac{d_{1}d_{2}}{(d_{1}-1)(d_{2}-1)}$. This means that the condition (9) on the maximum eigenvalue of $\Lambda^{T}\Lambda$ is satisfied. We then find that
$\operatorname{Tr}(\Lambda^{T}R)=\operatorname{Tr}(SL)=-\sqrt{\frac{d_{1}d_{2}}{(d_{1}-1)(d_{2}-1)}}\operatorname{Tr}(S).$
In this way we have obtained the lowest value of $\operatorname{Tr}(\Lambda^{T}R)$, while maintaining the condition (18) of block-positivity of $W$. Therefore, if this value is less than $-1$, which in view of (19) is equivalent to $\operatorname{Tr}(W\rho)<0$, we conclude that the state $\rho$ is entangled. In addition, one can show that $Tr(S)=Tr(\sqrt{RR^{T}})$. Therefore, we have the following simple criterion:
${\rm If\ \ }\gamma:=\operatorname{Tr}(S)>\sqrt{\frac{(d_{1}-1)(d_{2}-1)}{d_{1}d_{2}}},\ \ \ {\rm the\ state\ is\ entangled.}$ (20)
The condition that we derived does not require finding the elements of the matrix $\Lambda$, and only requires recasting the density matrix in terms of a basis chosen by orthogonal trace-less matrices, and calculating the singular values of the matrix $R$.
It should be mentioned that the singular values of $R$ and hence the value of $\gamma$ do not depend on the bases which we choose for the two parts for expanding the density matrix $\rho$, as long as these bases are orthonormal and traceless.
Remark: Interestingly, this criterion is a special case of Theorem 1 in [22], where the authors show that this is a powerful criterion and unifies several criteria known before.
We will now show by several examples that this criterion, originated from the EW defined in (12), is indeed powerful enough to detect entanglement of a class of PPT entangled states which has not been previously detected.
4 Examples
In this section, we consider a few examples and show the power of our criterion. In all the following examples, except example 4.2, the states are PPT entangled states reported in previous works. We will show that the criterion given in (20), can detect entanglement in a larger interval of the parameters previously reported, and it also can detect PPT entangled states with different dimensions, i.e. $d_{1}\neq d_{2}$. For checking the following examples, the reader should note that we have used the following order for Gellmann matrices:
[…]
$\textrm{for }\ \ \ \ \ \ \frac{1}{2}d(d-1)\leq i<d(d-1):\ \ \ \ \ \ \Gamma_{i}\in\Gamma_{y},$ (21)
$\textrm{for }\ \ \ \ \ \ d(d-1)\leq i<d^{2}:\hskip 28.45274pt\hskip 28.45274pt\ \ \Gamma_{i}\in\Gamma_{z},$ (22)
where for $1\leq m<n\leq d$,
$\Gamma_{x}=\big{\{}\ \ \frac{1}{\sqrt{2}}(|m\rangle\langle n|+|n\rangle\langle m|)\ \big{\}},$
$\Gamma_{y}=\big{\{}\ \ \frac{i}{\sqrt{2}}(|m\rangle\langle n|-|n\rangle\langle m|)\ \ \big{\}},$
and for $11\leq n\leq d-$ 1
$\Gamma_{z}=\big{\{}\frac{1}{\ \ \sqrt{n(n+1)}}(\sum_{m=0}^{n-1}|m\rangle\langle m|-n|n\rangle\langle n|)\ \ \big{\}}.$
Moreover in all those cases where the state is a PPT state and its partial transpose equals to itself, i.e. $\rho=\rho^{\Gamma}$, we can use the anti-symmetricity of the Gell-Mann matrices in $\Gamma_{y}$ to simplify the calculations in the following way:
Let at least one of the matrices $\Gamma_{i}$ or $\Omega_{j}$ (say $\Gamma_{i}$) be anti-symmetric, then we have
$r_{ij}=\operatorname{Tr}(\rho(\Gamma_{i}\otimes\Omega_{j}))=\operatorname{Tr}(\rho^{\Gamma}(\Gamma_{i}\otimes\Omega_{j}))=\operatorname{Tr}(\rho.(\Gamma_{i}^{T}\otimes\Omega_{j}))=-\operatorname{Tr}(\rho(\Gamma_{i}\otimes\Omega_{j}))=0.$
The same is true when $\Omega_{j}$ is anti-symmetric.
4.1 Example 1:
In this example, we demonstrate that the method explained in this paper can identify PPT entangled states, and thus, it can generate indecomposable EWs. It has been shown in [23] that the following state is PPT entangled state when $x\neq 1$ (for $x=1$ this is a separable state):
$\rho=\frac{1}{N}\left(\begin{array}[]{ccc|ccc|ccc}1&.&.&.&1&.&.&.&1\\ .&\frac{1}{x}&.&.&.&.&.&.&.\\ .&.&x&.&.&.&.&.&.\\ \hline\cr.&.&.&x&.&.&.&.&.\\ 1&.&.&.&1&.&.&.&1\\ .&.&.&.&.&\frac{1}{x}&.&.&.\\ \hline\cr.&.&.&.&.&.&\frac{1}{x}&.&.\\ .&.&.&.&.&.&.&x&.\\ 1&.&.&.&1&.&.&.&1\\ \end{array}\right).$ (23)
Here dots represent zeros, $N=3+x+x^{-1}$ is the normalization factor and $x>0$. The above state can be written in the basis of normalized Gell-Mann matrices as follows:
[…]
$+$ $\left(1-\frac{\xi}{2}\right)(\Gamma_{7}\otimes\Gamma_{7}+\Gamma_{8}\otimes\Gamma_{8})-\frac{\sqrt{3}}{2}\eta(\Gamma_{7}\otimes\Gamma_{8}-\Gamma_{8}\otimes\Gamma_{7})\Big{]},$
where $\xi=x+x^{-1}$ and $\eta=x-x^{-1}$. The eigenvalues of the matrix $R$ are easily found to be
$\lambda(R)=\frac{1}{N}\{1(3),-1(3),(1-\frac{\xi}{2})\pm i\frac{\sqrt{3}}{2}\eta\},$
where the numbers in parentheses indicate the multiplicity of each singular value. Thus, taking into account the normalization factor $N$, we find
$Tr(S)=\frac{6+2\sqrt{(1-\frac{\xi}{2})^{2}+\frac{3}{4}\eta^{2}}}{3(1+\xi)}.$
Condition (20) for entanglement of the state now amounts to
$\frac{1}{4}(\xi^{2}-\eta^{2})<\xi-1,$
or after simplification (inserting $\xi=x+x^{-1}$ and $\eta=x-x^{{}_{1}}$) leads to the following condition for the state (23) to be entangled:
$2<x+x^{-1}.$
This shows that the state $\rho_{x}$ is entangled for all values of $x$, except for $x=1$, which coincides with the result of [23].
4.2 Example 2:
This example demonstrates that our method can identify the entanglement of all NPT states in the following form in any dimension $d$:
$\rho=p|\psi\rangle\langle\psi|+(1-p)\frac{I}{d},\quad|\psi\rangle=\frac{1}{\sqrt{d}}\sum_{i=0}^{d-1}|i,i\rangle.$
To calculate $\gamma$ in equation (20), we need to obtain the coefficients $r_{ij}$:
$r_{ij}=p\operatorname{Tr}[\frac{1}{d}(\sum_{k,l}|kk\rangle\langle ll|).(\Gamma_{i}\otimes\Gamma_{j})]=\frac{p}{d}\operatorname{Tr}(\Gamma_{i}\Gamma_{j}^{T})=\pm\frac{p}{d}\delta_{ij}.$
Here we have noticed that the expression $\operatorname{Tr}(\Gamma_{i}\Gamma_{j}^{T})$ equals $\delta_{ij}$ when $\Gamma_{j}^{T}=\Gamma_{j}^{\dagger}$ and equals $-\delta_{ij}$ when $\Gamma_{j}^{\dagger}=-\Gamma_{j}^{T}$. As it is evident, this state is diagonal in the Gell-Mann matrix basis, and the absolute value of all its singular values is equal to $\frac{p}{d}$, and the number of its singular values is also equal to $d^{2}-1$. Therefore, we find for this state
$\gamma=\operatorname{Tr}(S)=\frac{p}{d}(d^{2}-1),$
which, according to eq (20), should be greater than $\frac{d-1}{d}$ so the state is entangled. Thus the range of $p$ for the state to be entangled is $p>\frac{1}{d+1}$. This is the same value obtained from the Peres criterion which in this case follows from
$\rho^{\Gamma}=(1-p)\frac{I}{d^{2}}+p\frac{1}{d}F,$
where $F$ is the swap operator. The eigenstates are the symmetric and antisymmetric states and the smallest eigenvalue, namely $\frac{1-p}{d^{2}}-\frac{p}{d}$, is negative when $p>\frac{1}{d+1}$ and hence the state is entangled. For this example the two methods give equivalent results. This result has already been reported in [24].
4.3 Example 3
In this example, we provide a PPT entangled state which can be detected by the criterion in equation (20). To construct a PPT entangled state we define the following Unextendible Product Basis in $C_{3}\otimes C_{4}:$
$|\phi_{1}\rangle=\frac{1}{\sqrt{2}}|0\rangle|1-2\rangle,\quad|\phi_{2}\rangle=\frac{1}{\sqrt{6}}|0\rangle(2|0\rangle-|1+2\rangle),\quad|\phi_{3}\rangle=\frac{1}{\sqrt{2}}|1\rangle|1-2\rangle,\quad$
$|\phi_{4}\rangle=\frac{1}{\sqrt{2}}|2\rangle|1-2\rangle,\quad|\phi_{5}\rangle=\frac{1}{\sqrt{6}}|2\rangle(|1+2\rangle-2|3\rangle),\quad|\phi_{6}\rangle=\frac{1}{\sqrt{2}}|1-2\rangle|0\rangle,\quad$ (25)
$|\phi_{7}\rangle=\frac{1}{\sqrt{2}}|0-1\rangle|3\rangle,\quad|\phi_{8}\rangle=\frac{1}{\sqrt{12}}|0+1+2\rangle|0+1+2+3\rangle,$ (26)
and construct the PPT entangled state $\rho=I-\sum_{i}|\phi_{i}\rangle\langle\phi_{i}|$ [8], whose explicit form is given by
$\rho=\frac{1}{48}\left(\begin{array}[]{c|c|c|c}3\ \epsilon&-\epsilon&-\epsilon&-\epsilon\\ \hline\cr-\epsilon&D&D^{\prime}&-\epsilon\\ \hline\cr-\epsilon&D^{\prime T}&D&-\epsilon\\ \hline\cr-\epsilon&-\epsilon&-\epsilon&3\ \epsilon\\ \end{array}\right),$
where $\epsilon$ is the 3 dimensional square matrix all of its elements are equal to 1, and
$D=\left(\begin{array}[]{ccc}5&-1&-1\\ -1&5&-1\\ -1&-1&5\end{array}\right),\quad D^{\prime}=\left(\begin{array}[]{ccc}-1&5&-1\\ -1&-1&5\\ 5&-1&-1\end{array}\right).$
The matrix $R$ is simply obtained by noting that its elements are given by $r_{mn}=\operatorname{Tr}(\rho(\Gamma_{m}\otimes\Omega_{n}))=-\sum_{i}\langle\phi_{i}|\Gamma_{m}\otimes\Omega_{n}|\phi_{i}\rangle$. The result is
$R=\frac{1}{24}\left(\begin{array}[]{c|c|c}-\epsilon_{3\times 6}&0_{3\times 6}&A_{3\times 3}\\ \hline\cr 0_{3\times 6}&0_{3\times 6}&0_{3\times 3}\\ \hline\cr B_{2\times 6}&0_{2\times 6}&C_{2\times 3}\end{array}\right),$
where $\epsilon_{3\times 6}$ is a $3\times 6$ matrix all of whose elements are 1, and similarly ${0}_{3\times 6}$ is a matrix all of whose elements are zero. Also
$A$ $=$ $\left(\begin{array}[]{ccc}0&0&-3\sqrt{\frac{3}{2}}\\ 0&0&0\\ 3&\sqrt{3}&\sqrt{\frac{3}{2}}\end{array}\right),$ (30)
$B$ $=$ $\left(\begin{array}[]{cccccc}{2}&{2}&-1&0&0&0\\ \frac{2}{\sqrt{3}}&\frac{2}{\sqrt{3}}&\frac{1}{\sqrt{3}}&0&\frac{-4}{\sqrt{3}}&\frac{-4}{\sqrt{3}}\end{array}\right),$ (33)
$C$ $=$ $\left(\begin{array}[]{ccc}0&0&-\frac{1}{2}\sqrt{\frac{3}{2}}\\ -\frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{5}{6\sqrt{2}}\end{array}\right).$ (36)
The singular values of this matrix are obtained numerically and one finds that
$\gamma=\operatorname{Tr}(S)=1.429>\sqrt{\frac{3\times 4}{2\times 3}}=\sqrt{2},$
which shows that while $\rho$ is a PPT entangled state, it can be detected by the criterion (20).
4.4 Example 4
In [23], among other things, the following family of PPT entangled states in $C_{4}\times C_{4}$ are introduced
$\rho=\frac{1}{N}\left(\begin{array}[]{cccc|cccc|cccc|cccc}1&.&.&.&.&1&.&.&.&.&1&.&.&.&.&1\\ .&\frac{1}{a}&.&.&.&.&.&.&.&.&.&.&.&.&1&.\\ .&.&1&.&.&.&.&.&1&.&.&.&.&.&.&.\\ .&.&.&a&.&.&1&.&.&.&.&.&.&.&.&.\\ \hline\cr.&.&.&.&a&.&.&.&.&.&.&1&.&.&.&.\\ 1&.&.&.&.&1&.&.&.&.&1&.&.&.&.&1\\ .&.&.&1&.&.&\frac{1}{a}&.&.&.&.&.&.&.&.&.\\ .&.&.&.&.&.&.&1&.&.&.&.&.&1&.&.\\ \hline\cr.&.&1&.&.&.&.&.&1&.&.&.&.&.&.&.\\ .&.&.&.&.&.&.&.&.&a&.&.&1&.&.&.\\ 1&.&.&.&.&1&.&.&.&.&1&.&.&.&.&1\\ .&.&.&.&1&.&.&.&.&.&.&\frac{1}{a}&.&.&.&.\\ \hline\cr.&.&.&.&.&.&.&.&.&1&.&.&\frac{1}{a}&.&.&.\\ .&.&.&.&.&.&.&1&.&.&.&.&.&1&.&.\\ .&1&.&.&.&.&.&.&.&.&.&.&.&.&a&.\\ 1&.&.&.&.&1&.&.&.&.&1&.&.&.&.&1\\ \end{array}\right),$ (37)
where $N=4a+\frac{4}{a}+8$, and it is shown that their entanglement can be detected by an entanglement witness, called $W_{ext}$, which shows that these states are entangled in the interval $10<a<$ 1. We now show that the interval of entanglement of these states is in fact much larger and covers the whole positive line $0<a<\infty$, except the point $a=1$. To show this, the $R$ matrix corresponding to (37) is calculated, which turns out to be
$R=\frac{1}{(a+1)^{2}}\big{(}A\oplus A^{\prime}\oplus B\big{)},$
where
$A=\left(\begin{array}[]{cccccc}\frac{a}{4}&.&.&.&.&\frac{a}{4}\\ .&\frac{a}{2}&.&.&.&.\\ .&.&\frac{a}{4}&\frac{a}{4}&.&.\\ .&.&\frac{a}{4}&\frac{a}{4}&.&.\\ .&.&.&.&\frac{a}{2}&.\\ \frac{a}{4}&.&.&.&.&\frac{a}{4}\end{array}\right),\ \ \ \ \ A^{\prime}=\left(\begin{array}[]{cccccc}-\frac{a}{4}&.&.&.&.&\frac{a}{4}\\ .&.&.&.&.&.\\ .&.&-\frac{a}{4}&-\frac{a}{4}&.&.\\ .&.&-\frac{a}{4}&-\frac{a}{4}&.&.\\ .&.&.&.&.&.\\ \frac{a}{4}&.&.&.&.&-\frac{a}{4}\end{array}\right),$
and
$B=\left(\begin{array}[]{ccc}-\frac{(a-1)^{2}}{8}&-\frac{a^{2}+2a-3}{8\sqrt{3}}&-\frac{(a-1)a}{2\sqrt{6}}\\ \frac{3a^{2}-2a-1}{8\sqrt{3}}&-\frac{(a-1)^{2}}{24}&-\frac{a^{2}+a-2}{6\sqrt{2}}\\ \frac{a-1}{2\sqrt{6}}&\frac{2a^{2}-a-1}{6\sqrt{2}}&-\frac{(a-1)^{2}}{12}\end{array}\right),$
with singular values given as
$\Big{\{}\frac{\left|a-1\right|}{4\left|a+1\right|}\ (2),\ \frac{\left|a\right|}{2(a+1)^{2}}\ (6),\ \frac{(a-1)^{2}}{4(a+1)^{2}}\ (1)\Big{\}}.$
This leads to
$\gamma=\frac{1}{4}\frac{(a-1)^{2}}{(a+1)^{2}}+3\frac{|a|}{(a+1)^{2}}+\frac{1}{2}\left|\frac{a-1}{a+1}\right|.$
Figure 1 shows that the value of $\gamma$ is always larger than $\sqrt{\frac{(d_{1}-1)(d_{2}-1)}{d_{1}d_{2}}}=\frac{3}{4}$ except for $a=1$.
Figure 1: Using the criterion intruduced in this paper, i.e. eq (17), we can see that state (37) is entangled for $0<a<\infty;a\neq 1$, which is much larger the range intruduced in [23].
Note that by $a<\infty$, we are emphasizing that $a$ should be a finite value. In fact, in the limit $a\longrightarrow\infty$, the state approaches to
$\rho_{\infty}=\frac{1}{4}\Big{(}|03\rangle\langle 03|+|10\rangle\langle 10|+|11\rangle\langle 11|+|22\rangle\langle 22|\Big{)},$
which is obviously separable. Incidentally, this is the limiting case where $\gamma$ approaches the value $\frac{3}{4}$.
4.5 Example 5
This example is a continuation of the previous example and demonstrates that for the second PPT entangled state introduced in [23], the interval of entanglement (indicated as $10<b<$ 1), is indeed much larger and as in the previous example covers all positive and finite values of $b$ different from $1$. The state is given by
$\rho=\frac{1}{N}\left(\begin{array}[]{cccc|cccc|cccc|cccc}1&.&.&.&.&1&.&.&.&.&1&.&.&.&.&1\\ .&\frac{1}{b}&.&.&1&.&.&.&.&.&.&.&.&.&.&.\\ .&.&1&.&.&.&.&.&1&.&.&.&.&.&.&.\\ .&.&.&b&.&.&.&.&.&.&.&.&1&.&.&.\\ \hline\cr.&1&.&.&b&.&.&.&.&.&.&.&.&.&.&.\\ 1&.&.&.&.&1&.&.&.&.&1&.&.&.&.&1\\ .&.&.&.&.&.&\frac{1}{b}&.&.&1&.&.&.&.&.&.\\ .&.&.&.&.&.&.&1&.&.&.&.&.&1&.&.\\ \hline\cr.&.&1&.&.&.&.&.&1&.&.&.&.&.&.&.\\ .&.&.&.&.&.&1&.&.&b&.&.&.&.&.&.\\ 1&.&.&.&.&1&.&.&.&.&1&.&.&.&.&1\\ .&.&.&.&.&.&.&.&.&.&.&\frac{1}{b}&.&.&1&.\\ \hline\cr.&.&.&1&.&.&.&.&.&.&.&.&\frac{1}{b}&.&.&.\\ .&.&.&.&.&.&.&1&.&.&.&.&.&1&.&.\\ .&.&.&.&.&.&.&.&.&.&.&1&.&.&b&.\\ 1&.&.&.&.&1&.&.&.&.&1&.&.&.&.&1\\ \end{array}\right),$
where $N=4b+\frac{4}{b}+8$, and its corresponding matrix is
$R=\frac{1}{(1+b)^{2}}\left(\frac{b}{2}I_{6}\oplus 0_{6}\oplus A\right),$
where $I_{6}$ and $0_{6}$ are respectively the identity matrix and the zero matrix in $C_{6}$ and
$A=\left(\begin{array}[]{ccc}-\frac{1}{8}(b-1)^{2}&-\frac{b^{2}+2b-3}{8\sqrt{3}}&-\frac{(b-1)b}{2\sqrt{6}}\\ \frac{3b^{2}-2b-1}{8\sqrt{3}}&-\frac{1}{24}(b-1)^{2}&-\frac{b^{2}+b-2}{6\sqrt{2}}\\ \frac{b-1}{2\sqrt{6}}&\frac{2b^{2}-b-1}{6\sqrt{2}}&-\frac{1}{12}(b-1)^{2}\\ \end{array}\right),$
with singular values given as
$\Big{\{}\frac{\left|b-1\right|}{4\left|b+1\right|}\ (2),\ \frac{\left|b\right|}{2(b+1)^{2}}\ (6),\ \frac{(b-1)^{2}}{4(b+1)^{2}}\ (1)\Big{\}},$
which is the same as the previous example and gives the same result.
4.6 Example 6
In [25], Terhal introduced a family of UPB for $3\times n$ systems. Here we take $n=4$ and show that our criterion can detect the PPT entangled state built by them in these dimensions.
The UPB states for the $3\times 4$ system are:
$|\psi_{0}\rangle$ $=$ $\frac{1}{\sqrt{2}}|0\rangle\otimes(|1\rangle-|3\rangle),$
$|\psi_{1}\rangle$ $=$ $\frac{1}{\sqrt{2}}|1\rangle\otimes(|2\rangle-|3\rangle),$ (38)
$|\psi_{2}\rangle$ $=$ $\frac{1}{\sqrt{2}}|2\rangle\otimes(|0\rangle-|3\rangle),$ (39)
$|\psi_{3}\rangle$ $=$ $\frac{1}{\sqrt{2}}(|0\rangle-|1\rangle)\otimes|0\rangle,$ (40)
$|\psi_{4}\rangle$ $=$ $\frac{1}{\sqrt{2}}(|1\rangle-|2\rangle)\otimes|1\rangle,$ (41)
$|\psi_{5}\rangle$ $=$ $\frac{1}{\sqrt{2}}(|2\rangle-|0\rangle)\otimes|2\rangle,$ (42)
$|\psi_{6}\rangle$ $=$ $\frac{1}{\sqrt{12}}\sum_{i=0}^{2}\sum_{j=0}^{3}|i\rangle\otimes|j\rangle.$ (43)
The PPT entangled state $\rho=I-\sum_{0}^{6}|\psi_{i}\rangle\langle\psi_{i}|$ turns out to be
$\rho=\frac{1}{60}\left(\begin{array}[]{ccc|ccc|ccc|ccc}5&-1&-1&-1&5&-1&-1&-1&-1&-1&-1&-1\\ -1&5&-1&5&-1&-1&-1&-1&-1&-1&-1&-1\\ -1&-1&5&-1&-1&-1&-1&-1&-1&-1&5&-1\\ \hline\cr-1&5&-1&5&-1&-1&-1&-1&-1&-1&-1&-1\\ 5&-1&-1&-1&5&-1&-1&-1&-1&-1&-1&-1\\ -1&-1&-1&-1&-1&5&-1&-1&-1&5&-1&-1\\ \hline\cr-1&-1&-1&-1&-1&-1&5&5&-1&-1&-1&-1\\ -1&-1&-1&-1&-1&-1&5&5&-1&-1&-1&-1\\ -1&-1&-1&-1&-1&-1&-1&-1&5&-1&-1&5\\ \hline\cr-1&-1&-1&-1&-1&5&-1&-1&-1&5&-1&-1\\ -1&-1&5&-1&-1&-1&-1&-1&-1&-1&5&-1\\ -1&-1&-1&-1&-1&-1&-1&-1&5&-1&-1&5\\ \end{array}\right).$ (56)
The corresponding $R$ matrix is given by
$R=\frac{1}{30}\left(\begin{array}[]{c|c|c}-\epsilon_{3\times 6}&0_{3\times 6}&A\\ \hline\cr 0_{3\times 6}&0_{3\times 6}&0_{3\times 3}\\ \hline\cr B&0_{2\times 6}&0_{2\times 3}\end{array}\right),$
where
$A=\left(\begin{array}[]{ccc}3&\sqrt{3}&\sqrt{\frac{3}{2}}\\ 0&-2\sqrt{3}&\sqrt{\frac{3}{2}}\\ -3&-\sqrt{3}&\sqrt{\frac{3}{2}}\end{array}\right),$
and
$B=\left(\begin{array}[]{cccccc}0&0&0&0&3&-3\\ 0&0&0&-2\sqrt{3}&\sqrt{3}&\sqrt{3}\\ \end{array}\right).$
The singular values of $R$ are found to be
$\left\{\frac{1}{2\sqrt{10}},\frac{1}{5\sqrt{2}},\frac{1}{5\sqrt{2}},\frac{1}{5\sqrt{2}},\frac{1}{5\sqrt{2}}\right\},$
and we find
$\gamma=\frac{1}{2\sqrt{10}}+\frac{4}{5\sqrt{2}}=\big{(}\frac{8+\sqrt{5}}{10}\big{)}\frac{1}{\sqrt{2}},$
which is greater than $\sqrt{\frac{(d_{1}-1)(d_{2}-1)}{d_{1}d_{2}}}=\frac{1}{\sqrt{2}}$. This shows that the PPT state (56) is entangled.
5 Conclusion
In this paper, a class of Entanglement Witnesses for two particle systems with different dimensions was presented. Specifically we showed in section (3), how the use of this class of EW’s can be reduced to a simple criterion, (equation (20)), which can be checked on the form of the density matrix. We then showed by checking several one-parameter family of PPT entangled states that this criterion can detect their entanglement in a region of parameters much larger than previously thought.