### v5

# Positive maps and Entanglement Witnesses in different dimensions

Abstract
We present a continuous, multiparameter family of positive maps between spaces of differing dimensions. This framework facilitates the construction of Entanglement Witnesses (EWs) specifically designed for systems in $d_{1}\times d_{2}$ dimensions. We derive a simple, closed-form criterion for detecting entanglement in general density matrices based on these witnesses. To demonstrate the effectiveness of this criterion, we apply it to a range of Positive Partial Transpose (PPT) entangled states, revealing that the parameter regions where these states exhibit entanglement are larger than previously reported. Furthermore, we prove that non-unital EWs, corresponding to non-unital maps, are not more powerful than unital EWs, thus supporting the focus on unital positive maps in recent studies. Our method complements existing approaches to separability criteria for density matrices in different dimensions.
Introduction
Any conceivable operation on a quantum system, be it a measurement or noisy evolution, can be represented by a Completely Positive Trace-preserving (CPT) map or a quantum channel for short. A map of this kind not only maps density matrices to density matrices, but all of its extensions also have this property. The name of completely positive map comes from this last property. At first sight it seems, due to this stringent completely positive property, that construction of CPT maps is much more difficult than that of simply positive maps. However it has turned out that it is quite the contrary. CPT maps can readily be constructed, among other methods, by resorting to their Kraus representation [1]. In fact any such map from a system $A$ to another system $B$ (possibly with different dimension can be written as
${\cal E}(\rho)=\sum_{\alpha=1}^{K}A_{\alpha}\rho A_{\alpha}^{\dagger}$ (1)
where $A_{\alpha}:H_{A}\longrightarrow H_{B}$ are arbitrary linear operators whose only condition is the trace-preserving property, $\sum_{\alpha=1}A_{\alpha}^{\dagger}A_{\alpha}=I_{A}$, a condition which can be easily fullfiled, i.e. one simply chooses arbitrary operators $A_{\alpha},\ \ \alpha=1,\cdots K-1$ and by multiplying them with suitable scale factors ensures that $\sum_{\alpha=1}^{K-1}A_{\alpha}^{\dagger}A_{\alpha}\leq I_{A}$ and defines the last Kraus operator simply as $A_{K}:=\sqrt{I-\sum_{\alpha=1}^{K-1}A_{\alpha}^{\dagger}A_{\alpha}}$. While construction of instances of CPT maps is not challenging, classification of these maps, specially in higher dimensions, and deciding whether or not a given map is CPT is generally difficult. On the other hand construction of instances of maps which are only positive, but not completely positive, is not straightforward at all and as expected, their classification has remained an open problem despite decades of efforts by physicists and mathematicians. The need for construction of instances of positive maps arises from their role in entanglement theory [2, 3, 4].
The identification of entangled states, in particular, presents a formidable challenge, underscored by its classification as an NP-Hard problem in mathematical terms [5]. Hence, the pursuit of viable techniques for detecting entangled states takes on paramount importance. Entangled states are referred to as states that are not separable, and separable states refer to states that can be written as follows[6]:
$\rho=\sum p_{i}\rho_{i}^{(1)}\otimes\rho_{i}^{(2)},$ (2)
where $0\leq p_{i}\leq 1$ and $\sum p_{i}=1$. It is not easy at all to determine whether or not a given state can be expressed in the aforementioned form. As a result, there is a pressing need for more effective methods for the detection of entanglement. Over time, diverse techniques have emerged for this purpose. The Peres criterion is considered one of the primary methods in this regard[7], i.e. a state which is separable satisifes
$(I\otimes T)(\rho)\geq 0,$ (3)
where $T$ is the transpose, hence the name partial transpose. Therefore, if the partial transpose of a quantum state is not positive, it must be entangled. However, this criterion alone cannot identify all types of quantum entangled states because there are entangled states known as PPT, whose partial transpose are positive. In fact separable states are a subset of PPT states and only when $d_{1}d_{2}\leq 6$, these two sets are equal. Here $d_{1}$ and $d_{2}$ are the dimensions of the two quantum systems. In higher dimensions one should search for other criteria of entanglement, many of which have been reviewed in [8]. The most important ones are called Entanglement Witnesses (EW), see [9] for a review. An EW refers to a Hermitian operator $W$, which has negative eigenvalues but has the following property:
$\operatorname{Tr}(W\sigma)\geq 0,\quad\forall\sigma\in S,$
where $S$ is the set of all separable states. Therefore, a negative expectation value for $W$ on a state $\rho$, then means that $\rho$ is entangled. Alternatively, one can think of $W$ as a block-positive operator, i.e.
$\langle\psi,\phi|W|\psi,\phi\rangle,\ \forall\ \psi\in H_{1}\ ,\forall\ \phi\in H_{2},$
which has negative eigenvalue when acting on the whole Hilbert space $H_{1}\otimes H_{2}$. Here comes the important link between entanlement witnesses and positive maps. In fact, it has been known that any entanglement witness $W$ can be written in the form
$W=(I\otimes\Phi)|\phi^{+}\rangle\langle\phi^{+}|,$
where $\Phi$ is a positive (but not completely positive) maps, and $|\phi^{+}\rangle=\frac{1}{\sqrt{d}}\sum_{\mu=0}^{d-1}|\mu\mu\rangle$, is a maximally entangled state[10]. In other words, $W$ is the Choi-matrix of the map $\Phi$, which due to its lack of complete positivity can have negative eigenvalue. Thus the problem of constructing entanglement witnesses becomes equivalent to constructing positive but not completely positive maps. This is an important and known result which connects the long-standing problem of classification of positive maps in pure mathematics to the new problem of detection of entanglement quantum information theory. Based on this equivalence, entanglement witnesses were introduced in many works, see [10, 9, 11, 8] and references therein. For more recent works in this direction, see [12, 13, 14], where the authors generally start from a restricted class of positive maps, mostly between spaces of equal dimensions and construct the corresponding EW’s. These maps are always uniital and their corresponding affine transformation is a rotation performed by an orthogonal matrix.
In this work, we construct a large class of positive maps which firstly act between spaces of different dimensions $d_{1}\times d_{2}$. Besides detecting entanglement in heterogenuous systems, this makes it possible to witness entanglement in a multi-partite system, for example when $k$ qubits are entnagled to $n-k$ qubits in a many-body system. In particular, we
a) provide sufficient conditions for our maps to be positive, but not completely positive, so that their Choi matrix is an EW, and then
b) provide sufficient conditions for our maps to be completely positive. Both criteria are expressed in terms of the affine transformation corresponding to this map.
c) We then prove that among the EW’s constructed from positive maps, the unital EW’s, that is, those constructed from unital maps are stronger than those corresponding to non-unital maps. Thus we rigorously support the focus on unital positive maps in recent studies [12, 13, 14].
d) We reduce the previous considerations to a simple criterion which solely depends on the density matrix itself and does not require any type of entanglement witnesses. Thus simply by checking this critria, one can decide whether or not a given state is entangled.
e) By studying several examples, we show that the EW’s generated in this way, can detect entanglement of states for a larger range of parameters than previously found.
f) Finally in an appendix, we restrict ourselves to unital EW’s and show how by using the concept of block-positivity, we can directly construct the EW’s derived from the positive maps.
The structure of this paper is as follows: In Section 1, we set up our notation. Then in section 2, we state Mehta’s lemma and its proof which is the basis of all our construction. In section 3, we use this lemma [15] to construct general positive maps between spaces of different dimensions. By general maps, we mean maps which are not necessarily unital. This part extends the previous works which have constructed EW’s along similar lines [16, 17, 18]. In section 4 we reduce the previous considerations to a simple criterion which solely depends on the density matrix itself and does not require any type of entanglement witnesses. We also show in this section that in each family of maps, the unital ones lead to the strongest type of witnesses. In Section 5, we present several examples which shows the power of our criterion in detecting PPT entangled states. We end up the paper with a conclusion. In an appendix, section 8, an alternative method based on the concept of block-positiviy is used to obtain the same results derived in section 3.
1 Notations and Conventions
Let $H_{d}$ be the Hilbert space of dimension $d$ and $L(H_{d})$ the space of linear operators acting on $H_{d}$. The half space of positive operators is denoted by $L(H_{d})^{+}$. We will be dealing with maps $\Phi:L^{+}(H_{d_{1}})\longrightarrow L^{+}(H_{d_{2}})$. The Hilbert-Schmidt norm of a real matrix $A$ is denoted by $\|A\|=\operatorname{Tr}(A^{T}A)^{1/2}$. By $S\in M_{d_{1}\times d_{2}}$, we mean a rectangular matrix of dimension $d_{1}\times d_{2}$, by $M\in O(d)$, we mean an orthogonal square matrix of dimension $d$. Let $A$ be a rectangular matrix. Then by $a_{max}$ we mean the largest eigenvalue of the positive symmetric matrix $\sqrt{A^{T}A}$ or equivalently $\sqrt{AA^{T}}$. A basis of orthonormal Hermitian and traceless matrices acting on $H_{d_{1}}$ is denoted by $\{\Gamma_{i},\ i=1\cdots d_{1}^{2}-1\}$ and the same basis of matrices acting on $H_{d_{2}}$ is denoted by $\{\Omega_{\mu},\mu=1,\cdots d_{2}^{2}-1\}$. Therefore we have $\operatorname{Tr}(\Gamma_{i})=\operatorname{Tr}(\Omega_{\mu})=0$ and $\operatorname{Tr}(\Gamma_{i}\Gamma_{j})=\delta_{ij}$ and $\operatorname{Tr}(\Omega_{\mu}\Omega_{\nu})=\delta_{\mu,\nu}.$ Finally we use the convention of summing over repeated indices. In all formulas, $i$ runs from $1$ to $d_{1}^{2}-1$ and $\mu$ from $1$ to $d_{2}^{2}-1$.
The state of a quantum system is denoted by $\rho\in D(H_{d})$, where $D(H_{d})\subset L(H_{d})^{+}$ is the convex subset of positive operators with unit trace. Let $X$ be a positive operator in $L^{+}(H_{d_{1}})$ which can be expanded as
$X={x_{0}}I_{d_{1}}+x_{i}\Gamma_{i}=x_{0}I_{d_{1}}+{\bf x}\cdot\bm{\Gamma},$ (4)
where $(x_{0},{\bf x}=(x_{1},\cdots x_{d_{1}^{2}-1}))$ are real parameters. For any positive operator $X$, it necessarily holds that $\operatorname{Tr}(X^{2})\leq\operatorname{Tr}(X)^{2}$ which leads to
${\bf x}\cdot{\bf x}\leq x_{0}^{2}d_{1}(d_{1}-1).$ (5)
where ${\bf x}\cdot{\bf x}:=x_{i}x_{i}$. If the operator $X$ is pure, that is, if $X=|\psi\rangle\langle\psi|$ for a non-normalized vector $|\psi\rangle$, then the equality sign holds in (5).
Any operator $Y$ in $L(H_{d_{2}})$ is written as
$Y={y_{0}}I_{d_{2}}+y_{\mu}\Omega_{\mu}=y_{0}I_{d_{2}}+{\bf y}\cdot\bm{\Omega}.$ (6)
Let $\Phi:D(H_{d_{1}})\longrightarrow D(H_{d_{2}})$ be a map which acts as follows:
$\Phi(I_{d_{1}})=(I_{d_{2}}+t_{\mu}\Omega_{\mu})\ \ \ \ ,\ \ \ \Phi(\Gamma_{i})=\Lambda_{\mu i}\Omega_{\mu},$ (7)
Here $t_{\mu}$ and $\Lambda_{\mu i}$ are real numbers which characterize the map $\Phi$. In the special case where ${\bf t}=0$, the map $\Phi$ is unital. Hereafter we use summation convention for summing over repeated indices. With these conventions, $\Phi$ maps $X$ to
$\Phi(X)={x_{0}}I_{d_{2}}+(x_{0}t_{\mu}+x_{i}\Lambda_{\mu i})\Omega_{\mu}=x_{0}I_{d_{2}}+({\bf t}x_{0}+\Lambda{\bf x})\cdot\Omega.$ (8)
With our convention, the map $\Phi$ corresponds to the following affine tranformation, $\Phi:(x_{0},{\bf x})\longrightarrow(x_{0},{\bf t}x_{0}+\Lambda{\bf x})$ which can be written in matrix form .
$\Phi:\begin{pmatrix}x_{0}\\ {\bf x}\end{pmatrix}\longrightarrow\begin{pmatrix}1&{\bf 0}\\ {\bf t}&\Lambda\end{pmatrix}\begin{pmatrix}x_{0}\\ {\bf x}\end{pmatrix}.$ (9)
where ${\bf t}$ is a $d_{2}$ dimensional vector and $\Lambda\in M_{d_{2}\times d_{1}}$.
2 Mehta’s Lamma
Let $A$ be a Hermitian matrix of dimension $D$, if
$\operatorname{Tr}(A^{2})\leq\frac{(\operatorname{Tr}A)^{2}}{D-1},$ (10)
then $A$ is positive.
Note that Mehta’s lemma provides only a sufficient condition. Its converse is not true, e.g. there are positive matrices which do not satisfy (10), as is readily seen from an example of a diagonal matrix $A=diag(1,2,x),\ \forall\ x\geq 6$.
Proof: The proof is taken from refernce [19]. Let $A$ be a Hermitian matrix of size $D$ and let $\alpha={\rm Tr}A/\sqrt{{\rm Tr}A^{2}}$. If $\alpha\geq\sqrt{D-1}$ then $A$ is positive. We calculate the traces in the eigenbasis of $A$. Let $(x_{1},\dots x_{D})$ denote the spectrum of $A$. Assume first that one eigenvalue, say $x_{1}$, is negative. Making use of the right hand side of the standard estimation between the ${l}_{1}$ and ${l}_{2}$–norms (with prefactor $1$) of a $D$–vector, $||A||_{2}\leq||A||_{1}\leq D||A||_{2}$, we infer
${\rm TrA}=\sum_{i=1}^{D}x_{i}<\sum_{i=2}^{D}x_{i}\leq\sqrt{D-1}\Bigl{(}\sum_{i=2}^{D}x_{i}^{2}\Bigr{)}^{1/2}<\sqrt{D-1}\sqrt{{\rm Tr}A^{2}}\ .$ (11)
This implies that $\alpha<\sqrt{D-1}$. Hence if the opposite is true and $\alpha\geq\sqrt{D-1}$ then none of the eigenvalues $x_{i}$ could be negative, so $A\geq 0$. $\square$.
3 Positive maps between spaces of different dimensions
Given a quantum map $\Phi$ characterized by the affine transformation on the generalized Bloch vector ${\bf r}\longrightarrow\Lambda{\bf r}+{\bf t}$, our goal is to find the condition on this affine map which gurantees positivity of $\Phi$. This is stated in the following:
3.1 Sufficient condition for positivity
Proposition: A map $\Phi$, implementing the affine transformatoin (9) is positive if the following condition holds:
$\|{\bf t}\|+\sqrt{d_{1}(d_{1}-1)}\lambda_{max}\leq\sqrt{\frac{d_{2}}{d_{2}-1}},$ (12)
where $\lambda_{max}$ is the largest eigenvalue of the positive matrix $\sqrt{\Lambda\Lambda^{T}}.$
Proof: Due to linearity of the map, it suffices to prove the statement only for pure states. For such states (see (5)), we have
$\operatorname{Tr}(\Phi(X))=x_{0}d_{2}$
and
$\operatorname{Tr}(\Phi(X)^{2})=x_{0}^{2}d_{2}+\|{\bf t}x_{0}+\Lambda{\bf x}\|^{2}={x_{0}^{2}}d_{2}+x_{0}^{2}{\bf t}\cdot{\bf t}+2x_{0}{\bf t}\cdot\Lambda{\bf x}+\|\Lambda{\bf x}\|^{2}.$
In view of the Mehta lemma, the map $\Phi$ is positive if for all $(x_{0},{\bf x})$ subject to the condition (10), the following inequality holds
${x_{0}^{2}}d_{2}+x_{0}^{2}{\bf t}\cdot{\bf t}+2x_{0}{\bf t}\cdot\Lambda{\bf x}+\|\Lambda{\bf x}\|^{2}\leq\frac{x_{0}^{2}d_{2}^{2}}{d_{2}-1}.$ (13)
We demand a slightly stronger condition, (at a price of loosing some of the positive maps). To this end, we note that
$\|\Lambda{\bf x}\|^{2}\leq\lambda_{max}^{2}{\bf x}\cdot{\bf x}=\lambda_{max}^{2}x_{0}^{2}d_{1}(d_{1}-1),$ (14)
where $\lambda_{max}^{2}$ is the largest eigenvalue of the positive matrix $\Lambda^{T}\Lambda$. We also use Cauchy-Schwartz inequality to write
${\bf t}\cdot\Lambda{\bf x}\leq\|{\bf t}\|\ \|\Lambda{\bf x}\|\leq\|{\bf t}\|\lambda_{max}\|{\bf x}\|=\|{\bf t}\|\lambda_{max}{x_{0}}\sqrt{d_{1}(d_{1}-1)},$ (15)
Inserting (14) and (15) in (13) and dividing both sides by $x_{0}^{2}$, we find the following sufficient condition for positivity of the map $\Phi$
$\|{\bf t}\|^{2}+2\lambda_{max}\|{\bf t}\|\sqrt{d_{1}(d_{1}-1)}+\lambda_{max}^{2}d_{1}(d_{1}-1)\leq\frac{d_{2}}{(d_{2}-1)}.$ (16)
Taking the square root of both sides leads to (12) which concludes the proof. $\square$
For unital maps where ${\bf t}=0$, the above condition simplifies to
$\lambda_{max}\leq\sqrt{\frac{d_{2}}{d_{1}(d_{1}-1)(d_{2}-1)}}.$
3.2 Sufficient condition for complete positivity
A map $\Phi$ is completely positive if all its extensions are also positive. We now derive a sufficient condition on the affine map for complete positivity of the corresponding map.
Proposition: A positive map (8), implementing the affine transformatoin (9) is a completely positive map if the following condition holds:
${\bf t}\cdot{\bf t}+d_{1}\operatorname{Tr}(\Lambda\Lambda^{T})\leq\frac{d_{2}}{d_{1}d_{2}-1}.$ (17)
Proof: According to a celebrated theorem of Choi [20], a map is completely positive if its Choi matirx defined as below is positive
$C_{\Phi}$ $=$ $(I_{d_{1}}\otimes\Phi)(|\phi^{+}\rangle\langle\phi^{+}|)=(|i\rangle\langle j|\otimes\Phi(|i\rangle\langle j|)$ (18)
$=$ $|i\rangle\langle j|\otimes\Big{[}\frac{1}{d_{1}}\delta_{ij}(I_{d_{2}}+{\bf t}\cdot\Omega)+\langle j|\Gamma_{l}|i\rangle\Lambda_{\mu l}\Omega_{\mu}\Big{]}$ (19)
$=$ $\frac{1}{d_{1}}I_{d_{1}}\otimes(I_{d_{2}}+{\bf t}\cdot\Omega)+\Lambda_{\mu l}\Gamma_{l}^{T}\otimes\Omega_{\mu}.$ (20)
from which the traces are easily derived as
$\operatorname{Tr}(C_{\Phi})=d_{2},\hskip 28.45274pt\operatorname{Tr}(C_{\Phi}^{2})=\frac{d_{2}+{\bf t}\cdot{\bf t}}{d_{1}}+\operatorname{Tr}(\Lambda\Lambda^{T})$ (21)
Inserting these in Mehta’s Lemma (10), leads to (17) and completes the proof $\square$.
Example: Maps between qubits: As an example consider the qubit maps $\Phi$ with the affine transformation (9) where $\Lambda=\begin{pmatrix}\lambda_{1}&&\\ &\lambda_{2}&\\ &&\lambda_{3}\end{pmatrix}$ and ${\bf t}=(t_{1},t_{2},t_{3})^{T}$. Then the condition for positivity of the map is given from (12) to be
$\sqrt{t_{1}^{2}+t_{2}^{2}+t_{3}^{2}}+\sqrt{2}\lambda_{max}\leq\sqrt{2},$ (22)
and the condition of complete positivity is given from (17) to be
$\sum_{i}(t_{i}^{2}+2\lambda_{i}^{2})\leq\frac{2}{3}.$ (23)
We shoud emphasize that since the Mehta’s condition is a sufficient conditions, the above inequalities determine only a subset of positive and completely positive maps. For example, if we restrict ourselves to unital maps, then the condition (17) is given by
$\lambda_{1}^{2}+\lambda_{2}^{2}+\lambda_{3}^{2}\leq\frac{1}{3},$
which is a sphere of radius $r=\frac{1}{\sqrt{3}}$ and is a subset of the Fujiwara-Algoet tetrahedron
$|\lambda_{1}+\lambda_{2}|\leq 1+\lambda_{3}\ \ \ \ \ |\lambda_{1}-\lambda_{2}|\leq 1-\lambda_{3}.$
A comparison of the volumes of these two sets shows that $30$ percents of the completely positive unital maps for qubits are detected by this simple use of Mehta’s criteria, figure (1).
Figure 1: Example of qubit maps, with eigenvalues of affine transformation given by $\lambda_{1},\lambda_{2}$ and $\lambda_{3}$. Restricting to unital maps, the tethdrahedron shows the complete set of unital maps as derived from the Fujiwara-Algoet condition, the sphere shows those maps which are derived from Mehta’s condition. Its volume is approximately 30 percent of the volume of the tethrahedron.
4 Entanglement Witnesses from positive maps
It is known that there is a one-to-one correspondence between positive map which are not completely positive on the one hand and Entanglement Witnesses on the other. The EW corresponding to the map (8) is given by [10]
$W=(I_{d_{1}}\otimes\Phi)|\beta\rangle\langle\beta|,$
where $|\beta\rangle=\frac{1}{\sqrt{d_{1}}}\sum_{i=0}^{d_{1}-1}|i,i\rangle$ is a maximally entangled state. Direct calculation, after multiplication by an overall factor $d_{1}d_{2}$, leads to the following witness
$W=I_{d_{1}}\otimes I_{d_{2}}+t_{\mu}I_{d_{1}}\otimes\Omega_{\mu}+d_{1}{\Lambda}_{\mu i}\Gamma_{i}\otimes\Omega_{\mu}.$ (24)
In the last step, we have used the freedom to replace the matrices $\Gamma_{i}^{T}$ by $\Gamma_{i}$, since both sets can be chosen as orthonormal set for the space of Hermitian traceless matrices on $H_{d_{1}}$. Given a general bi-partite density matrix as
$\rho=\frac{1}{d_{1}d_{2}}I_{d_{1}}\otimes I_{d_{2}}+{a}_{i}\Gamma_{i}\otimes I_{d_{2}}+{b}_{\mu}I_{d_{1}}\otimes\Omega_{\mu}+{r}_{i\mu}\Gamma_{i}\otimes\Omega_{\mu},$ (25)
one finds that
$\operatorname{Tr}(W\rho)=1+d_{1}{\bf b}\cdot{\bf t}+d_{1}\operatorname{Tr}(R\Lambda).$ (26)
Those EW’s which correspond to unital positive maps (i.e. ${\bf t}=0$) are called unital entanglement witness. Therefore for any two dimensions, any operator of the above form can be an entanglement witness, provided that the condition (12) holds. Whether or not an EW constructed in this way, has also a negative eigenvalue depends on the matrix $\Lambda$ that we choose. In all previous papers that were mentioned in the introduction, $\Lambda$ should be orthogonal. These comprise a much smaller class compared to the $\Lambda$ matrices in this paper. Moreover, finding orthogonal matrices in high dimensions is not a straightforward task. In this paper, one can use any matrix with any properties to construct entanglement witnesses, as long as eq. (12) is satisfied. For any given density matrix $\rho$ characterized by the pair $({\bf b},R)$ as shown above, we have to find a witness characterized by a pair $({\bf t},\Lambda)$ such that $\operatorname{Tr}(W\rho)$ as written above is minimized. Note that for this kind of witness the parameter ${\bf a}$ in the density matrix plays no role. This minimization should be done in the range of all possible entanglement witnesses, namely among all those affine transormations which satisfy the inequality (65), which we rewrite here for convenience:
$\|{\bf t}\|+\sqrt{d_{1}(d_{1}-1)}\lambda_{max}\leq\sqrt{\frac{d_{2}}{d_{2}-1}}.$ (27)
To proceed, we need the following:
Definiiton: A unital entanglement witness is one which corresponds to the Choi matrix of a unital positive map.
We now show that show that for any given state $\rho$, if a non-unital EW can detect its entanglement, there is a unital EW whcih can detect it in a more effective way. This provides a rigorous justification for the focus of researchers in the past years on trying to construct unital positive maps and their corresponding entanglement witnesses.
Theorem: Given a general bi-partites state $\rho$, unital entanglement witnesses (e.g. those derived from unital maps) are more effective in detecting entanglement of this state.
Proof: In $\operatorname{Tr}(W\rho)$ given in (26) the vector ${\bf b}$ and the matrix $R$ are determined by the density matrix $\rho$ and our task is to choose the vector ${\bf t}$ and the matrix $\Lambda$ to make $\operatorname{Tr}(W\rho)$ as negative as possible. To do so, we invoke the Chauchy Schwartz inequality and choose ${\bf t}$ anti-parallel to ${\bf b}$, that is we take ${\bf t}=-\alpha{\bf b}$, where $\alpha$ is positive constant which can be taken as large as possible provided that the condition (12), namely
$\|{\bf t}\|+\sqrt{d_{1}(d_{1}-1)}\lambda_{max}\leq\sqrt{\frac{d_{2}}{d_{2}-1}}$
is not violated.
To minimize $\operatorname{Tr}(\Lambda R)$, let us use the singular value decomposition of $R$, i.e. $R=USV$, where $U,V$ are rotation matrices respectively belonging to orthogonal groups $O(d_{1})$ and $O(d_{2})$ and $S\in M_{d_{1}\times d_{2}}$ is a rectangular matrix on whose main diagonal are the singular values of $R$. Then using the freedom of choosing $\Lambda$, we write it as $\Lambda=V^{T}LU^{T}$, where we should choose $L$ such that $\operatorname{Tr}(\Lambda R)$ is minimized subject to the condition (12). Thus we have
$\operatorname{Tr}(\Lambda R)=\operatorname{Tr}(LS).$ (28)
Noting that the singular values of $R$ (i.e. the diagonal entries of $S$) are all positive, this equation can be minimized by considering $L$ to have on its diagonal, a non-negative parameter denoted by $-\beta$ and all the other off-diagonal elements to be zero. As explained in detail in appendix I, this leads to $\operatorname{Tr}(R\Lambda)=-\beta\|R\|_{1}.$ We then find
$\operatorname{Tr}(W\rho)=1-\alpha\|{\bf b}\|^{2}-d_{1}\beta\|R\|_{1}.$ (29)
We have to find the minimum of this quantity in the region $\alpha-\beta$ defined by
$\alpha\|{\bf b}\|+\sqrt{d_{1}(d_{1}-1)}\beta\leq\sqrt{\frac{d_{2}}{d_{2}-1}}.$ (30)
This is now a simple mathematical problem which can be expressed most simply when we cast it as follows: We want to find the minimum value of the function $F(\alpha,\beta):=\operatorname{Tr}(\rho W)$
$F(\alpha,\beta):=1-p\alpha-q\beta,$
in the region
$p^{\prime}\alpha+q^{\prime}\beta\leq 1,$
where
$p=\|{\bf b}\|^{2},\ \ \ q=d_{1}\|R\|_{1},\ \ \ p^{\prime}={\|{\bf b}\|}\sqrt{\frac{d_{2}-1}{d_{2}}},\ \ \ q^{\prime}=\sqrt{\frac{d_{1}(d_{1}-1)(d_{2}-1)}{d_{2}}}.$ (31)
The function $F$ being a linear function of $\alpha$ and $\beta$ aquires its minimum on one of the three vertices of the triangle, figure (2).
Figure 2: We should minimize the amount of $\operatorname{Tr}(W\rho)=1-p\alpha-q\beta$ with respect to the gray area, where $p,q,p^{\prime},q^{\prime}$ are difined in eq. (31).
On the vertex $(0,0)$, we have $F(0,0)=1,$ which is positive, hence no detection of entanglement. To explore the value of $F$ on the other two vertices, we calculate the value of $F$ on the oblique edge of the triangle defined by $\alpha p^{\prime}+\beta q^{\prime}=1$. On this edge, the function $F$ depends on one parameter $\alpha$, which after eliminating $\beta$ in favor of $\alpha$, allows us to define the following one-parameter function $\tilde{F}(\alpha)=\operatorname{Tr}(\rho W)$
$\tilde{F}(\alpha)=1-p\alpha-q(\frac{1-\alpha p^{\prime}}{q^{\prime}})=1-\frac{q}{q^{\prime}}-\alpha p^{\prime}(\frac{p}{p^{\prime}}-\frac{q}{q^{\prime}}).$ (32)
Note that the parameters $p,p^{\prime},q$ and $q^{\prime}$ come from the density matrix $\rho$ and $\alpha$ comes from the witness $W$ which is to be chosen so that the value of $\tilde{F}$ is as negative as possible. In case that $\frac{p}{p^{\prime}}<\frac{q}{q^{\prime}}$, we take $\alpha$ to be zero leading to the minimum value
$\tilde{F}_{min}=\tilde{F}(\alpha=0)=1-\frac{q}{q^{\prime}}.$ (33)
In case that $\frac{p}{p^{\prime}}>\frac{q}{q^{\prime}}$, we take $\alpha$ to be as large as possible, namely $\frac{1}{p^{\prime}}$ which gives the minimum value
$\tilde{F}_{min}=\tilde{F}(\alpha=\frac{1}{p^{\prime}})=1-\frac{p}{p^{\prime}}.$ (34)
But the condition $\frac{p}{p^{\prime}}<\frac{q}{q^{\prime}}$ already shows that $\tilde{F}(\alpha=0)\leq\tilde{F}(\alpha=\frac{1}{p^{\prime}})$ which means that there is no advantage in using non-unital witnesses, that is, any entangled state which can be detected by a non-unital witness ($\alpha=\frac{1}{q^{\prime}}$), is also detected by a unital witness ($\alpha=0$). Those states which are detected by this unital witness are those for which $\tilde{F}(0)=1-\frac{q^{\prime}}{q}<0$. Inserting the values of $q^{\prime}$ and $q$ from (31), this leads to the following condition:
${\rm If\ \ }\|R\|_{1}>\sqrt{\frac{(d_{1}-1)(d_{2}-1)}{d_{1}d_{2}}},\ \ \ \ \ \ {\rm the\ state\ is\ entangled.}$ (35)
We stress the significance of this result which is now independent of any entanglemet witnesses and is a criterion based entirely on the density matrix itself. We will now show by several examples that this criterion, originated from the EW defined in (24), is indeed powerful enough to detect entanglement of a class of PPT entangled states and indeed in some examples, the range of parameters where the state is shown to be entangled is much larger than previously reported.
5 Examples
In this section, we consider a few examples and show the power of our criterion. This criterion which is derived here by starting from Mehta’s lemma and positive maps, is a special case of a more general criteria already reported in [21]. In all the following examples, except example 5.1, the states are PPT entangled states reported in previous works [22, 23]. We will show that the criterion given in (35), can detect entanglement in a larger interval of the parameters previously reported, and it also can detect PPT entangled states with different dimensions, i.e. $d_{1}\neq d_{2}$. For checking the following examples, the reader should note that we have used the following order for Gellmann matrices:
[…]
$\textrm{for }\ \ \ \ \ \ \frac{1}{2}d(d-1)\leq i<d(d-1):\ \ \ \ \ \ \ \ \ \Gamma_{i}\in\Gamma_{y},$ (36)
$\textrm{for }\ \ \ \ \ \ d(d-1)\leq i<d^{2}:\hskip 28.45274pt\hskip 28.45274pt\ \ \Gamma_{i}\in\Gamma_{z},$ (37)
where for $1\leq m<n\leq d$,
$\Gamma_{x}=\big{\{}\ \ \frac{1}{\sqrt{2}}(|m\rangle\langle n|+|n\rangle\langle m|)\ \big{\}},$
$\Gamma_{y}=\big{\{}\ \ \frac{i}{\sqrt{2}}(|m\rangle\langle n|-|n\rangle\langle m|)\ \ \big{\}},$
and for $11\leq n\leq d-$ 1
$\Gamma_{z}=\big{\{}\frac{1}{\ \ \sqrt{n(n+1)}}(\sum_{m=0}^{n-1}|m\rangle\langle m|-n|n\rangle\langle n|)\ \ \big{\}}.$
Moreover in all those cases where the state is a PPT state and its partial transpose equals to itself, i.e. $\rho=\rho^{\Gamma}$, we can use the anti-symmetricity of the Gell-Mann matrices in $\Gamma_{y}$ to simplify the calculations in the following way:
Let at least one of the matrices $\Gamma_{i}$ or $\Omega_{j}$ (say $\Gamma_{i}$) be anti-symmetric, then we have
$r_{ij}=\operatorname{Tr}(\rho(\Gamma_{i}\otimes\Omega_{j}))=\operatorname{Tr}(\rho^{\Gamma}(\Gamma_{i}\otimes\Omega_{j}))=\operatorname{Tr}(\rho.(\Gamma_{i}^{T}\otimes\Omega_{j}))=-\operatorname{Tr}(\rho(\Gamma_{i}\otimes\Omega_{j}))=0.$
The same is true when $\Omega_{j}$ is anti-symmetric.
5.1 Example 1:
This example demonstrates that our method can identify the entanglement of all NPT states in the following form in any dimension $d$:
$\rho=p|\psi\rangle\langle\psi|+(1-p)\frac{I}{d},\quad|\psi\rangle=\frac{1}{\sqrt{d}}\sum_{i=0}^{d-1}|i,i\rangle.$
It is readily found that
$r_{ij}=p\operatorname{Tr}[\frac{1}{d}(\sum_{k,l}|kk\rangle\langle ll|).(\Gamma_{i}\otimes\Gamma_{j})]=\frac{p}{d}\operatorname{Tr}(\Gamma_{i}\Gamma_{j}^{T})=\pm\frac{p}{d}\delta_{ij}.$
Here we have noticed that the expression $\operatorname{Tr}(\Gamma_{i}\Gamma_{j}^{T})$ equals $\delta_{ij}$ when $\Gamma_{j}^{T}=\Gamma_{j}^{\dagger}$ and equals $-\delta_{ij}$ when $\Gamma_{j}^{\dagger}=-\Gamma_{j}^{T}$. As it is evident, this state is diagonal in the Gell-Mann matrix basis, and the absolute value of all its singular values is equal to $\frac{p}{d}$, and the number of its singular values is also equal to $d^{2}-1$. Therefore, we find for this state
$\|R\|_{1}=\operatorname{Tr}(\sqrt{R^{T}R})=\frac{p}{d}(d^{2}-1),$
which, according to eq (35), should be greater than $\frac{d-1}{d}$ for the state to be entangled. Thus the range of $p$ for the state to be entangled is $p>\frac{1}{d+1}$. This is the same value obtained from the Peres criterion which in this case follows from
$\rho^{\Gamma}=(1-p)\frac{I}{d^{2}}+p\frac{1}{d}F,$
where $F$ is the swap operator. The eigenstates are the symmetric and antisymmetric states and the smallest eigenvalue, namely $\frac{1-p}{d^{2}}-\frac{p}{d}$, is negative when $p>\frac{1}{d+1}$ and hence the state is entangled. For this example the two methods give equivalent results. This result has already been reported in [24].
5.2 Example 2
In [22], among other things, the following family of PPT entangled states in $C_{4}\times C_{4}$ are introduced
$\rho=\frac{1}{N}\left(\begin{array}[]{cccc|cccc|cccc|cccc}1&.&.&.&.&1&.&.&.&.&1&.&.&.&.&1\\ .&\frac{1}{a}&.&.&.&.&.&.&.&.&.&.&.&.&1&.\\ .&.&1&.&.&.&.&.&1&.&.&.&.&.&.&.\\ .&.&.&a&.&.&1&.&.&.&.&.&.&.&.&.\\ \hline\cr.&.&.&.&a&.&.&.&.&.&.&1&.&.&.&.\\ 1&.&.&.&.&1&.&.&.&.&1&.&.&.&.&1\\ .&.&.&1&.&.&\frac{1}{a}&.&.&.&.&.&.&.&.&.\\ .&.&.&.&.&.&.&1&.&.&.&.&.&1&.&.\\ \hline\cr.&.&1&.&.&.&.&.&1&.&.&.&.&.&.&.\\ .&.&.&.&.&.&.&.&.&a&.&.&1&.&.&.\\ 1&.&.&.&.&1&.&.&.&.&1&.&.&.&.&1\\ .&.&.&.&1&.&.&.&.&.&.&\frac{1}{a}&.&.&.&.\\ \hline\cr.&.&.&.&.&.&.&.&.&1&.&.&\frac{1}{a}&.&.&.\\ .&.&.&.&.&.&.&1&.&.&.&.&.&1&.&.\\ .&1&.&.&.&.&.&.&.&.&.&.&.&.&a&.\\ 1&.&.&.&.&1&.&.&.&.&1&.&.&.&.&1\\ \end{array}\right),$ (38)
where $N=4a+\frac{4}{a}+8$, and it is shown that their entanglement can be detected by an entanglement witness, called $W_{ext}$, which shows that these states are entangled in the interval $10<a<$ 1. We now show that the interval of entanglement of these states is in fact much larger and covers the whole positive line $0<a<\infty$, except the point $a=1$. To show this, the $R$ matrix corresponding to (38) is calculated, which turns out to be
$R=\frac{1}{(a+1)^{2}}\big{(}A\oplus A^{\prime}\oplus B\big{)},$
where
$A=\left(\begin{array}[]{cccccc}\frac{a}{4}&.&.&.&.&\frac{a}{4}\\ .&\frac{a}{2}&.&.&.&.\\ .&.&\frac{a}{4}&\frac{a}{4}&.&.\\ .&.&\frac{a}{4}&\frac{a}{4}&.&.\\ .&.&.&.&\frac{a}{2}&.\\ \frac{a}{4}&.&.&.&.&\frac{a}{4}\end{array}\right),\ \ \ \ \ A^{\prime}=\left(\begin{array}[]{cccccc}-\frac{a}{4}&.&.&.&.&\frac{a}{4}\\ .&.&.&.&.&.\\ .&.&-\frac{a}{4}&-\frac{a}{4}&.&.\\ .&.&-\frac{a}{4}&-\frac{a}{4}&.&.\\ .&.&.&.&.&.\\ \frac{a}{4}&.&.&.&.&-\frac{a}{4}\end{array}\right),$
and
$B=\left(\begin{array}[]{ccc}-\frac{(a-1)^{2}}{8}&-\frac{a^{2}+2a-3}{8\sqrt{3}}&-\frac{(a-1)a}{2\sqrt{6}}\\ \frac{3a^{2}-2a-1}{8\sqrt{3}}&-\frac{(a-1)^{2}}{24}&-\frac{a^{2}+a-2}{6\sqrt{2}}\\ \frac{a-1}{2\sqrt{6}}&\frac{2a^{2}-a-1}{6\sqrt{2}}&-\frac{(a-1)^{2}}{12}\end{array}\right),$
with singular values given as
$\Big{\{}\frac{\left|a-1\right|}{4\left|a+1\right|}\ (2),\ \frac{\left|a\right|}{2(a+1)^{2}}\ (6),\ \frac{(a-1)^{2}}{4(a+1)^{2}}\ (1)\Big{\}}.$
This leads to
$\|R\|_{1}=\operatorname{Tr}(\sqrt{R^{T}R})=\frac{1}{4}\frac{(a-1)^{2}}{(a+1)^{2}}+3\frac{|a|}{(a+1)^{2}}+\frac{1}{2}\left|\frac{a-1}{a+1}\right|.$
Figure 3 shows that the value of $\gamma$ is always larger than $\sqrt{\frac{(d_{1}-1)(d_{2}-1)}{d_{1}d_{2}}}=\frac{3}{4}$ except for $a=1$.
Figure 3: Using the criterion intruduced in this paper, i.e. eq (65), we can see that state (38) is entangled for $0<a<\infty;a\neq 1$, which is much larger the range intruduced in [22].
Note that by $a<\infty$, we are emphasizing that $a$ should be a finite value. In fact, in the limit $a\longrightarrow\infty$, the state approaches to
$\rho_{\infty}=\frac{1}{4}\Big{(}|03\rangle\langle 03|+|10\rangle\langle 10|+|11\rangle\langle 11|+|22\rangle\langle 22|\Big{)},$
which is obviously separable. Incidentally, this is the limiting case where $\gamma$ approaches the value $\frac{3}{4}$.
5.3 Example 3
In [23], Terhal introduced a family of UPB for $3\times n$ systems. Here we take $n=4$ and show that our criterion can detect the PPT entangled state built by them in these dimensions.
The UPB states for the $3\times 4$ system are:
$|\psi_{0}\rangle$ $=$ $\frac{1}{\sqrt{2}}|0\rangle\otimes(|1\rangle-|3\rangle),$
$|\psi_{1}\rangle$ $=$ $\frac{1}{\sqrt{2}}|1\rangle\otimes(|2\rangle-|3\rangle),$ (39)
$|\psi_{2}\rangle$ $=$ $\frac{1}{\sqrt{2}}|2\rangle\otimes(|0\rangle-|3\rangle),$ (40)
$|\psi_{3}\rangle$ $=$ $\frac{1}{\sqrt{2}}(|0\rangle-|1\rangle)\otimes|0\rangle,$ (41)
$|\psi_{4}\rangle$ $=$ $\frac{1}{\sqrt{2}}(|1\rangle-|2\rangle)\otimes|1\rangle,$ (42)
$|\psi_{5}\rangle$ $=$ $\frac{1}{\sqrt{2}}(|2\rangle-|0\rangle)\otimes|2\rangle,$ (43)
$|\psi_{6}\rangle$ $=$ $\frac{1}{\sqrt{12}}\sum_{i=0}^{2}\sum_{j=0}^{3}|i\rangle\otimes|j\rangle.$ (44)
The PPT entangled state $\rho=I-\sum_{0}^{6}|\psi_{i}\rangle\langle\psi_{i}|$ turns out to be
$\rho=\frac{1}{60}\left(\begin{array}[]{ccc|ccc|ccc|ccc}5&-1&-1&-1&5&-1&-1&-1&-1&-1&-1&-1\\ -1&5&-1&5&-1&-1&-1&-1&-1&-1&-1&-1\\ -1&-1&5&-1&-1&-1&-1&-1&-1&-1&5&-1\\ \hline\cr-1&5&-1&5&-1&-1&-1&-1&-1&-1&-1&-1\\ 5&-1&-1&-1&5&-1&-1&-1&-1&-1&-1&-1\\ -1&-1&-1&-1&-1&5&-1&-1&-1&5&-1&-1\\ \hline\cr-1&-1&-1&-1&-1&-1&5&5&-1&-1&-1&-1\\ -1&-1&-1&-1&-1&-1&5&5&-1&-1&-1&-1\\ -1&-1&-1&-1&-1&-1&-1&-1&5&-1&-1&5\\ \hline\cr-1&-1&-1&-1&-1&5&-1&-1&-1&5&-1&-1\\ -1&-1&5&-1&-1&-1&-1&-1&-1&-1&5&-1\\ -1&-1&-1&-1&-1&-1&-1&-1&5&-1&-1&5\\ \end{array}\right).$ (57)
The corresponding $R$ matrix is given by
$R=\frac{1}{30}\left(\begin{array}[]{c|c|c}-\epsilon_{3\times 6}&0_{3\times 6}&A\\ \hline\cr 0_{3\times 6}&0_{3\times 6}&0_{3\times 3}\\ \hline\cr B&0_{2\times 6}&0_{2\times 3}\end{array}\right),$
where
$A=\left(\begin{array}[]{ccc}3&\sqrt{3}&\sqrt{\frac{3}{2}}\\ 0&-2\sqrt{3}&\sqrt{\frac{3}{2}}\\ -3&-\sqrt{3}&\sqrt{\frac{3}{2}}\end{array}\right),$
and
$B=\left(\begin{array}[]{cccccc}0&0&0&0&3&-3\\ 0&0&0&-2\sqrt{3}&\sqrt{3}&\sqrt{3}\\ \end{array}\right).$
The singular values of $R$ are found to be
$\left\{\frac{1}{2\sqrt{10}},\frac{1}{5\sqrt{2}},\frac{1}{5\sqrt{2}},\frac{1}{5\sqrt{2}},\frac{1}{5\sqrt{2}}\right\},$
and we find
$\|R\|_{1}=\operatorname{Tr}(\sqrt{R^{T}R})=\frac{1}{2\sqrt{10}}+\frac{4}{5\sqrt{2}}=\big{(}\frac{8+\sqrt{5}}{10}\big{)}\frac{1}{\sqrt{2}},$
which is greater than $\sqrt{\frac{(d_{1}-1)(d_{2}-1)}{d_{1}d_{2}}}=\frac{1}{\sqrt{2}}$. This shows that the PPT state (57) is entangled.
6 Conclusion
Mehta’s lemma is a simple sufficient condition for a matrix to be positive. We have used this lemma to construct a large family of positive maps and characterized them in terms of their corresponding affine transformations, i.e. transformation of the so-called Bloch vector ${\bf r}\longrightarrow\Lambda{\bf r}+{\bf t}$, equation (12). Using these positive maps, and by using the one-to-one correspondence between entanglement witnesses and the Choi matrix of positive maps, a multi-parameter family of Entanglement Witnesses for two particle systems with different dimensions was constructed. Specifically we showed how the use of this class of EW’s can be reduced to a simple criterion, (equation (35)), which can be checked on the form of the density matrix. We then showed by checking several one-parameter family of PPT entangled states that this criterion can detect their entanglement in a region of parameters much larger than previously thought. We also rigorously proved that unital entanglement witnesses, i.e. those corresponding to unital positive maps, can detect a larger set of entangled states.
Acknowledgement: This research was supported in part by Iran National Science Foundation, under Grant No.4022322. The authors wish to thank S. Roofeh, A. Farmanian, and A. Tangestani, for many inspiring discussions during this project. The authors also thank Dariusz Chruscinski for his valuable comments.